A seed company mixes two types of seed for bird feeding. One costs $1.10 per kg and the other costs $2.25 per kg. How much of each type of seed is needed to produce 6kg at a cost of $8.90?
Answer by checkley77(12569)

2.25x+1.10(6-x)=8.90
2.25x+6.60-1.10x=8.90
1.15x=8.90-6.60
1.15x=2.30
x=2.30/1.15
x=2 kg of the $2.25 seed is used.
6-2=4 kg of the $1.10 seed is used.
Proof:
2.25*2+1.10*4=8.9
4.50+4.40=8.90
8.90=8.90